17. Gummo Corporation just paid a dividend of $0.75 per share, and that dividend is expected to grow at a constant rate of 6.5% per year. The company's beta is 0.95, the required return on the market is 10.5%, and the risk-free rate is 4.5%. What is the company's current stock price?

a) 10.02

b)20.27

c)9.40

d)21.59

e)16.98
Answer #1

required return= risk-free rate +Beta*(market rate- risk-free rate )

=4.5+0.95*(10.5-4.5)

=10.2%

Current price=D1/(Required return-Growth rate)

=(0.75*1.065)/(0.102-0.065)

which is equal to

=$21.59(Approx).